OpenStudy (anonymous):

Janice has 8 DVD case on a shelf, one for each season of her favorite TV show .Her brother accidentally knocks them off the shelf onto the floor .When her brother puts them back on the shelf ,he does not pay attention to the season numbers and puts the case back on the shelf randomly .Find each probability 1)P(season 5 in the correct position) 2)P(season 1 and 8 in the correct position)

OpenStudy (anonymous):

1)\[\frac{1}{8}\]

OpenStudy (anonymous):

2) \[\frac{1}{8}\times \frac{1}{7}\]

OpenStudy (anonymous):

can you help me more?

OpenStudy (anonymous):

P(season 1 through 4 in the correct position)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

P(all even number seasons followed by all odd-numbered seasons)

OpenStudy (anonymous):

we reason as follows: probability season one is in correct spot is 1/8 then the probability season two is in the correct spot GIVEN season one is is 1/7

OpenStudy (anonymous):

P(all even number seasons in the correct position)

OpenStudy (anonymous):

P(season 5 through 8 in ant order followed by seasons 1 through 4 in any oder)

OpenStudy (anonymous):

and so on. then multiply to get \[\frac{1}{8}\times \frac{1}{7}\times \frac{1}{6}\times \frac{1}{5}\]

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

that was answer to P(season 1 through 4 in the correct position)

OpenStudy (anonymous):

P(all even number seasons followed by all odd-numbered seasons) ?

OpenStudy (anonymous):

all even numbers in correct position. only one way to do that, but then 4! ways to arrange the other ones so that is \[\frac{4!}{8!}=\frac{1}{8\times 7\times 6\times 5}\]

OpenStudy (anonymous):

P(season 5 through 8 in ant order followed by seasons 1 through 4 in any oder) \[\frac{4!\times 4!}{8!}\]

OpenStudy (anonymous):

hope it is clear how i am getting these

OpenStudy (anonymous):

all even numbers followed by all odd numbers \[\frac{4}{8}\times \frac{3}{7}\times \frac{2}{6}\times \frac{1}{5}\] and once you have all the even numbers first the odds must follow

OpenStudy (anonymous):

got this?